PET has a Young's, Modulus of 3.5 GPa and PLGA 1.2 GPa. A polymer blend is prepared with 40:60 mixture (PET:PlGA) . What is the Young's, Modulus of this blend (use mixture rule). Calculate the percent load borne by PET when the blend is subjected to a load. If after 10 days, half of PET and one third of PLGA had degraded, what is the Young's Modulus of this blend.


1
Expert's answer
2021-04-12T06:58:44-0400

Explanations & Calculations

  • When mixed, the property of elastic modulus varies according to the mixing ratio.
  • Then that is for the blend will be,

"\qquad\qquad
\begin{aligned}
\small Y_{blend}&=\small 3.5\,GPa\times \frac{40}{100}+1.2\,GPa\times \frac{60}{100}\\
&=\small \bold{2.12\,GPa}
\end{aligned}"

"\qquad\qquad ..............................................................."


  • When a load is applied on the material it behaves according to the properties of each component materials. A rough estimate of the percent load carried by PET can be obtained by comparing the moduli of the blend & the particular component material.
  • Then,

"\qquad\qquad
\begin{aligned}
\small \frac{F_{PET}}{F_{blend}}\times100&=\small \frac{3.5\times 0.4}{2.12}\times 100=\bold{66.04}
\end{aligned}"

"\qquad\qquad..............................................................."


  • After the degradation, there exist obsolete material within the blend which causes reduction in properties.
  • To calculate the new modulus, the percent amounts of effective material is needed.
  • The total volume of the blended material does not change. If the initial volume of PET was "\small V_1" and that of PIGA was "\small V_2", then at the beginning

"\qquad\qquad
\begin{aligned}
\small V_{pet}:V_{piga}&=\small 40:60=2:3
\end{aligned}"

  • What is left effective after 10 days is

"\qquad\qquad
\begin{aligned}
\small V_{pet}&=\frac{2}{5}\times \frac{1}{2}=\frac{1}{5}
\\
\small V_{PIGA}&=\small \frac{3}{5}\times \frac{2}{3}=\frac{2}{5}
\end{aligned}"

  • Then their ratio is

"\qquad\qquad
\begin{aligned}
\small V_{PET}:V_{PIGA}:V_{obsolete\,material}
&=\small 100\times \frac{1}{5}:100\times \frac{2}{5}:100\times\frac{2}{5}\\
&=\small 20:40:40
\end{aligned}"

  • Then the new modulus will be

"\qquad\qquad
\begin{aligned}
\small Y'&=\small 3.5\,GPa\times\frac{20}{100}+2.1\,GPa\times\frac{40}{100}\\
&=\small 0.7+0.84\\
&=\small \bold{1.54 \,GPa}
\end{aligned}"